OpenStudy (anonymous):

inverse laplace transform of (s^2-3s+1)/(s^3+s^2-2s)

OpenStudy (anonymous):

the denominator will be s(s+2)(s-1), but im not sure what to do with the numerator?

hartnn (hartnn):

tried partial fraction ?

OpenStudy (anonymous):

does it split into 3 fractions though? im really trying to remember partial fractions its been so long

hartnn (hartnn):

yes, it does,... \(\dfrac{s^2-3s+1}{s(s+2)(s-1)}=\dfrac{A}{s}+\dfrac{B}{s+2}+\dfrac{C}{s-1}=\)

hartnn (hartnn):

know how to find A,B,C ?

OpenStudy (anonymous):

umm well to find A you would divide s^2+s-2 into s^2-3s+1?? that doesnt seem right though

hartnn (hartnn):

first, do this, s^2-3s+1 = A(....) + B (.....) +C (.....) can you fill the blanks ? simple algebra

OpenStudy (anonymous):

actually i dont remember this at all, are you talking about just putting the s^2 into the first, and -3s into the second and 1 into the third??

hartnn (hartnn):

no...i was asking you to make a common denominator for left side, like w/x+y/z = (wz+xy)/xz

hartnn (hartnn):

FOR RIGHT SIDE***

hartnn (hartnn):

\(\dfrac{s^2-3s+1}{s(s+2)(s-1)}=\dfrac{A}{s}+\dfrac{B}{s+2}+\dfrac{C}{s-1} \\ \implies s^2-3s+1= A (s+2)(s-1)+Bs(s-1)+Cs(s+2) \) did you get this ???

OpenStudy (anonymous):

ohhh yeah, thats what i was saying with s^2+s-2... for A, I guess since you didn't say anything I was confused... but yes, im totally with you.

hartnn (hartnn):

to find A, put s=0 what u get ?

OpenStudy (anonymous):

i just noticed that i typed it in wrong... it is s^2-3s-1. but to find A, are u saying plug in 0 for s and you get -2A and you solve setting that equal to s^2-3s-1... so then just divide by -2

OpenStudy (anonymous):

wait then it would be 1/2

hartnn (hartnn):

ok, then \(s^2-3s-1= A (s+2)(s-1)+Bs(s-1)+Cs(s+2)\) A= 0 -1 =A (-2) yes, A =1/2 correct, now you know how can you find B ?

OpenStudy (anonymous):

well first, why did you pick s=0?? you can't use that for B now though

hartnn (hartnn):

look at this equation closely \(s^2-3s-1= A (s+2)(s-1)+Bs(s-1)+Cs(s+2)\) this is true for every value of 's' now if i want to find A , what value of 's' can i put so that i will not get B and C in calculations ?

OpenStudy (anonymous):

gotcha, so to find b we use -2 and to find c we use 1.

hartnn (hartnn):

absolutely correct!

OpenStudy (anonymous):

k, i will get those.

OpenStudy (anonymous):

B=3/2 and C=1

OpenStudy (anonymous):

c=-1

hartnn (hartnn):

correct :) now take the inverse LT of individual terms, \(\dfrac{s^2-3s+1}{s(s+2)(s-1)}=\dfrac{1}{2s}+\dfrac{3}{2(s+2)}-\dfrac{1}{s-1}\) can you ?

OpenStudy (anonymous):

well then f(x) = 1/2 + 3/2e^-2x - e^-x *just followed the table of inverse LT.

hartnn (hartnn):

yes, i get the same..... i would have used variable 't' instead..

OpenStudy (anonymous):

is there any particular reason why use t instead?

hartnn (hartnn):

habit.... if F (x) = (....) then use 'x' if F (t) = (....) then use 't' else use any variable...